Let f(x) = (1 + x2),1. Find the radius of convergence of the Taylor series of f about x, = 0. fix 2Let f(x) = (1 + x2),1. Find the radius of convergence of the Taylor series of f about x, = 0. fix 2
Answer #1

solution. Given Junction f/8)= (1+x2) U- Taylor series of of ť about to +24x + f(x)= 1-x² {(-) x ²n neo 2n comparaing with a